OpenStudy (anonymous):

Which is a factor of 18x^2+ 3x – 6 ? 2x + 1 2x + 3 3x + 2 3x – 1

OpenStudy (kira_yamato):

18x^2 + 3x – 6 = 3 (6x^2 + x - 2) = 3 (2x + 1) (3x - 2) 2x + 1 is a factor